# ul Koszykowa 75, 00-662 Warsaw, Poland

### v4

@@ Abstract.
We expand a hypergeometric function in an orthogonal series of Jacobi polynomials. We first identify identities involving the Pochhammer symbol (rising factorial). We utilize them to discover closed forms for certain integrals of Jacobi polynomials that are multiplied by a hypergeometric function and a Beta density. We can also obtain closed forms for particular series that consist of rising factorials, which generalize binomial series, by using well-known properties of the hypergeometric function. We can also get some simplifying identities of generalized hypergeometric functions.
@@
Pochhammer symbol, Hypergeometric function, Beta distribution, Jacobi polynomials
@@ 1. Introduction and Notation
As stated in the abstract, the paper will expand a hypergeometric function in an orthogonal series of Jacobi polynomials and find some related Beta integrals of the hypergeometric function. We will briefly recall basic notions used in the sequel to complete the paper. Then we will obtain specific, non-trivial, useful identities involving the Pochhammer symbol and related to Jacobi polynomials. This is done in Section 2. We will use the obtained identities in Section 3 to get several expansions of the hypergeometric function in the orthogonal series of Jacobi polynomials. We can also get some simplifying identities of generalized hypergeometric functions, as it is done by the end of Section 3.
@@
with $\left(a\right)_{\left(0\right)}\allowbreak=\allowbreak 1$. Let us notice that we have
$\left(a\right)^{\left(n\right)}\allowbreak=\allowbreak\left(-1\right)^{n}\left(-a\right)_{\left(n\right)}$
and of course
$\left(a\right)_{\left(n\right)}\allowbreak=\allowbreak\left(-1\right)^{n}\left(-a\right)^{\left(n\right)}.$
@@
It is commonly known that (1.2), can be derived from the formula below, known to be a particular case of the Gauss theorem giving the value of $_{2}F_{1}\left(a;b,c;1\right)$ (defined by (3.1, below) for $a\allowbreak=\allowbreak-n:$
@@ 2. Pochhammer symbol and Jacobi polynomials
3) For all complex $x,y$ whose sum is not equal to an integer less or equal to $1$ and $n\geq 0,$ we have
(2.4) $1=\sum_{s=0}^{n}\binom{n}{s}\frac{\left(x+s\right)^{\left(n-s\right)}\left(y\right)^{\left(s\right)}}{\left(y+x+s-1\right)^{\left(s\right)}\left(y+x+2s\right)^{\left(n-s\right)}}.$
@@
Let us recall the definition of Beta distribution. On one hand, we have the distribution with the density:
@@
It is also known that
@@
One can also show, by the simple change of the variable under the integral that
@@
i) The first of this identities is shown, e.g., in (see [4] (4.14)). The second follows the fact that if $x\in[0,1],$ then $2x-1\in[-1,1]$ and that $2h(2x-1|a,b)\allowbreak=\allowbreak f(x|a,b).$ So by the suitable change of variables orthogonality of polynomials $\left\{J_{n}\right\}$ with respect to $h$ is equivalent to orthogonality of polynomials $\left\{K_{n}\right\}$ with respect to $f$.
Now let us prove the third one. We have
@@
iii) The value at 11 is obvious, following (2.11). For the value at 0 we have again following (2.11),
$K_{n}(0|a,b)$ $=\frac{\left(-1\right)^{n}}{n!}\sum_{m=0}^{n}(-1)^{n-m}\binom{n}{m}\left(a+b+n-1\right)^{\left(m\right)}(b+m)^{(n-m)}$
$=\frac{\left(-1\right)^{n}}{n!}\left(a\right)^{\left(n\right)}.$
In the last line we applied formula (2.9) of [6]. ∎
As an immediate corollary, we get the following identities.
@@
(2.25) $\frac{1}{a+b-1}\sum_{j=0}^{n}\left(-1\right)^{j}\binom{n}{j}\left(a+b+j\right)^{\left(n-1\right)}$ $=\delta_{n,0},$
@@
Again, Tom Koornwinder, in a private communication, noticed that (2.21) can be obtained from the so-called Pfaff-Saalschütz formula ( http://dlmf.nist.gov/16.4.E3) applied to the generalized hypergeometric function $_{3}F_{2}\left(\begin{array}[c]{ccc}-n&a+c+n-1&b\\ c&a+b&\end{array};1\right).$ For the definition of hypergeometric function, see the section 3.
@@ 3. Hypergeometric function
Now, the third object of the paper is the so-called hypergeometric function, studied already by L. Euler. It is a function defined by the following infinite series
@@
considered for all complex $\left|x\right|<1,$ such that $c$ is not a non-positive integer. We will also use the simplified form of the hypergeometric function, namely $_{2}F_{1}\left(a,b;c;x\right)$. If $\operatorname{Re}(c-a-b)>0$ then one extends the domain of $x$ by adding the unit circle , i.e., when $\left|x\right|\allowbreak=\allowbreak 1$. To avoid unnecessary complications and also take into account that we intend to expand the hypergeometric function in the orthogonal series, we will always assume that parameters $a,$ $b,$ $c$ are real and such that cab>0c-a-b>0. After obtaining the orthogonal expansion, we can extend it from the real segment $[0,1]$ to the unit circle on the complex plane.
Let us also notice that
$_{2}F_{1}\left(a,b;a;x\right)=(1-x)^{-b}.$
@@
(3.7) $\frac{\left(b\right)^{\left(n\right)}}{n!}_{2}F_{1}\left(\begin{array}[c]{cc}-n&a+b+n-1\\ b&\end{array};x\right)$ $=K_{n}(1-x|a,b),$
(3.10) $\frac{\left(a\right)^{\left(n\right)}}{n!}~F_{1}\left(\begin{array}[c]{cc}-n&a+b+n-1\\ a&\end{array};x\right)$ $=(-1)^{n}K_{n}(x|a,b).$
@@
Let us notice that the notion of hypergeometric function can be easily generalized to the following function (also sometimes called hypergeometric )
$_{n}F_{m}\left(\begin{array}[c]{ccc}a_{1}&\ldots&a_{n}\\ b_{1}&\ldots&b_{m}\end{array};x\right)=\sum_{j\geq 0}\frac{\prod_{k=1}^{n}\left(a_{k}\right)^{\left(j\right)}}{\prod_{k=1}^{m}\left(b_{k}\right)^{\left(j\right)}}\frac{x^{j}}{j!}.$
Let us start with the following result.
Lemma 2.
@@
(3.14) $_{3}F_{2}\left(\begin{array}[c]{ccc}a+j&b+j&f+j\\ c+j&d+f+2j&\end{array};1\right),$
@@
(3.18) $\times~_{3}F_{2}\left(\begin{array}[c]{ccc}a+j&b+j&f+j\\ c+j&d+f+2j&\end{array};1\right).$
@@
(3.19) $_{2}F_{1}\left(a,b;c;x\right)=\left(c-1\right)\sum_{j\geq 0}\frac{(-1)^{j}\left(a\right)^{\left(j\right)}\left(b\right)^{\left(j\right)}x^{j}}{\left(c-1+j\right)j!}~_{2}F_{1}\left(a+j,b+j;j+1;x\right)$
(3.20) $=(1-x)^{1-a-b}\left(c-1\right)$

### v5

@@ Abstract.
We first present some identities involving the Pochhammer symbol (rising factorial). We also recall and present some new properties of the Jacobi polynomials. We use them to expand a general hypergeometric function in an orthogonal series of Jacobi polynomials. Then we use these expansions to discover closed forms for certain integrals of Jacobi polynomials that are multiplied by a generalized hypergeometric function and a Beta density. We can also obtain closed forms for some series involving rising factorials that generalise binomial series by using well-known properties of the hypergeometric function. In particular, we get a few new, nontrivial identities involving the Pochhammer symbol. We can also derive some simplifying identities for generalized hypergeometric functions.
@@
Pochhammer symbol, hypergeometric function, generalized hypergeometric function, Beta distribution, Jacobi polynomials
@@ 1. Introduction and Notation
As stated in the abstract, the paper will expand a generalized hypergeometric function in an orthogonal series of Jacobi polynomials and find some related Beta integrals of the generalized hypergeometric function. As a byproduct, we will obtain a few nontrivial identities involving the Pochhammer symbol (the rising factorial).
Before proceeding with this expansion, we will briefly recall the basic notions used in the sequel to complete the paper. Then we will obtain specific, nontrivial, useful identities involving the Pochhammer symbol and related to Jacobi polynomials. This is done in Section 2. We will use the obtained identities in Section 3 to get several expansions of the hypergeometric function in the orthogonal series of Jacobi polynomials. We can also derive some simplifying identities for hypergeometric functions, as discussed at the end of Section 3.
@@
with $\left(a\right)_{\left(0\right)}\allowbreak=\allowbreak 1$.
Let us notice that we have
$\left(a\right)^{\left(n\right)}\allowbreak$ $=\allowbreak\left(-1\right)^{n}\left(-a\right)_{\left(n\right)},$
$\left(a\right)_{\left(n\right)}\allowbreak$ $=\allowbreak\left(-1\right)^{n}\left(-a\right)^{\left(n\right)},$
and of course direcly from the definition we have
(1.2) $\left(-x-n+1\right)^{\left(n\right)}=\left(-1\right)^{n}\left(x\right)^{\left(n\right)}.$
@@
It is commonly known that (1.3), can be derived from the formula below, known to be a particular case of the Gauss theorem giving the value of $_{2}F_{1}\left(a;b,c;1\right)$ (defined by (3.2, below) for $a\allowbreak=\allowbreak-n:$
@@ 2. Pochhammer symbol and Jacobi polynomials
3) For all complex $x,y$ whose sum is not equal to an integer less or equal to 11 and $n\geq 0,$ we have
(2.4) $1=\sum_{j=0}^{n}\binom{n}{j}\frac{\left(x+j\right)^{\left(n-j\right)}\left(y\right)^{\left(j\right)}}{\left(y+x+j-1\right)^{\left(j\right)}\left(y+x+2j\right)^{\left(n-j\right)}}.$
@@
Let us recall the definition of Beta distribution.
On one hand, we have the distribution with the density:
@@
Remark 2.
Let’s make it clear that polynomials $\left\{J_{n}\right\}$ and $\left\{K_{n}\right\}$ are defined for all complex values of parameters $a$ and $b$, but they are orthogonal with respect to positive measure only when they are actually positive.
@@
It is also known that for $n\geq 1$ we have
@@
One can also show, by the simple change of the variable under the integral that for $n\geq 1$ we have
@@
i) The first of these identities is shown, e.g., in (see [5] (4.14)). The second one follows the fact that if $x\in[0,1],$ then $2x-1\in[-1,1]$ and that $2h(2x-1|a,b)\allowbreak=\allowbreak f(x|a,b).$ So by a suitable change of variables orthogonality of polynomials $\left\{J_{n}\right\}$ with respect to $h$ is equivalent to orthogonality of polynomials $\left\{K_{n}\right\}$ with respect to $f$.
Now, let us prove the third one. We have
@@
iii) The value at $1$ is obvious, following (2.11). For the value at 0 we have apply (2.18). ∎
The following identities are received immediately as a corollary.
@@
(2.25) $\left(a+b-1\right)\sum_{j=0}^{n}\left(-1\right)^{j}\binom{n}{j}\left(a+b+j\right)^{\left(n-1\right)}$ $=\delta_{n,0},$
@@
Again, Tom Koornwinder, in a private communication, noticed that (2.21) can be obtained from the so-called Pfaff-Saalschütz formula
( http://dlmf.nist.gov/16.4.E3) applied to the generalized hypergeometric function $_{3}F_{2}\left(\begin{array}[c]{ccc}-n&a+c+n-1&b\\ c&a+b&\end{array};1\right).$ For the definition of hypergeometric function, see the section 3.
@@ 3. Hypergeometric function
Now, the third object analysed in the paper is the so-called generalized hypergeometric function defined by the following infinite series
(3.1) $_{n}F_{m}\left(\begin{array}[c]{ccc}a_{1}&\ldots&a_{n}\\ b_{1}&\ldots&b_{m}\end{array};x\right)=\sum_{j\geq 0}\frac{\prod_{k=1}^{n}\left(a_{k}\right)^{\left(j\right)}}{\prod_{k=1}^{m}\left(b_{k}\right)^{\left(j\right)}}\frac{x^{j}}{j!}.$
Although one can formally consider this function for all non-negative integers $n$ and $m,$ we will consider it for all non-negative integers $n\leq m+1.$ This is so because for $n>m+1$ there appear some convergence problems of the defining series (3.1). For details, see [4] and [6]. Following [6], 16.2(iii) we will assume that when $n\leq m$ then all $b_{1},\ldots,b_{m}$ have positive real parts positive reals, when $n=m+1,$ then we will assume that $\operatorname{Re}(\sum_{j=1}^{m}b_{j}-\sum_{j=1}^{n}a_{j})>0$ and of course we assume that $\left|x\right|\leq 1$. In general let us notice that whenever $\prod_{k=1}^{n}a_{k}\allowbreak=\allowbreak 0,$ then $_{n}F_{m}\left(\begin{array}[c]{ccc}a_{1}&\ldots&a_{n}\\ b_{1}&\ldots&b_{m}\end{array};x\right)\allowbreak=\allowbreak 1$ for all $x.$ Also when one of the parameters $a_{1},\ldots,a_{n}$ is equal to a negative integer, say equal to n-n, then $_{n}F_{m}\left(\begin{array}[c]{ccc}a_{1}&\ldots&a_{n}\\ b_{1}&\ldots&b_{m}\end{array};x\right)$ is a polynomial in $x$ of order $n$.
Let us remark that the case $n\allowbreak=\allowbreak 2$ and $m\allowbreak=\allowbreak 1$ is treated specially. The function $_{2}F_{1}\left(\begin{array}[c]{cc}a&b\\ c&\end{array};x\right)$ is called the hypergeometric function, It was studied already by L. Euler, but its many major properties were obtained by Gauss. It is will be alternatively denoted as $_{2}F_{1}\left(a,b;c;x\right).$ Hence, we $_{2}F_{1}$ is the function defined by the following infinite series
@@
considered for all complex $\left|x\right|<1,$ such that $c$ is not a non-positive integer. If $\operatorname{Re}(c-a-b)>0$ then one extends the domain of $x$ by adding the unit circle , i.e., when $\left|x\right|\allowbreak=\allowbreak 1$. To avoid unnecessary complications and also take into account that we intend to expand the hypergeometric function in the orthogonal series, we will always assume that parameters $a,$ $b,$ $c$ are real and such that cab>0c-a-b>0. After obtaining the orthogonal expansion, we can extend it from the real segment $[0,1]$ to the unit circle on the complex plane.
Let us also notice that for all positive $m,a_{1},\ldots,a_{m}$ we have
$_{m+1}F_{m}\left(\begin{array}[c]{cccc}a_{1}&\ldots&a_{m}&b\\ a_{1}&\ldots&a_{m}&\end{array};x\right)=(1-x)^{-b}.$
@@
(3.8) $\frac{\left(b\right)^{\left(n\right)}}{n!}~_{2}F_{1}\left(\begin{array}[c]{cc}-n&a+b+n-1\\ b&\end{array};x\right)$ $=K_{n}(1-x|a,b),$
(3.11) $\frac{\left(a\right)^{\left(n\right)}}{n!}~_{2}F_{1}\left(\begin{array}[c]{cc}-n&a+b+n-1\\ a&\end{array};x\right)$ $=(-1)^{n}K_{n}(x|a,b).$
@@
Let us start with the following result. $\operatorname{Re}$
Theorem 1.
For any complex $a_{1},\ldots,a_{n},$ $\operatorname{Re}(b_{1})>0,\ldots\operatorname{Re}(b_{m})>0$ and $\operatorname{Re}(\sum_{j=1}^{m}b_{j}-\sum_{j=1}^{n}a_{j})>0$ and $\left|x\right|\leq 1$ and positive reals $f,d$ we have:
i)
(3.14) $_{n}F_{m}\left(\begin{array}[c]{ccc}a_{1}&\ldots&a_{n}\\ b_{1}&\ldots&b_{m}\end{array};(1+x)/2\right)=$
$\sum_{k\geq 0}J_{k}(x|f,d)\frac{\prod_{s=1}^{n}\left(a_{s}\right)^{\left(k\right)}}{\prod_{s=1}^{m}\left(b_{s}\right)^{\left(k\right)}\left(d+f+k-1\right)^{\left(k\right)}}$
(3.17) $\times~_{n+1}F_{m+1}\left(\begin{array}[c]{cccc}a_{1}+k&\ldots&a_{n}+k&f+k\\ b_{1}+k&\ldots&b_{m}+k&\left(d+f+2k\right)\end{array};1\right)$
ii)
(3.20) $_{n}F_{m}\left(\begin{array}[c]{ccc}a_{1}&\ldots&a_{n}\\ b_{1}&\ldots&b_{m}\end{array};x\right)=$
$\sum_{j\geq 0}K_{j}(x|f,d)\frac{\prod_{s=1}^{n}\left(a_{s}\right)^{\left(j\right)}}{\prod_{s=1}^{m}\left(b_{s}\right)^{\left(j\right)}\left(d+f+j-1\right)^{\left(j\right)}}$
(3.23) $\times~_{n+1}F_{m+1}\left(\begin{array}[c]{cccc}a_{1}+j&\ldots&a_{n}+j&f+j\\ b_{1}+j&\ldots&b_{m}+j&\left(d+f+2j\right)\end{array};1\right).$
iii) Assume additionally that $b_{1}\allowbreak\neq$$1,$ then
(3.26) $_{n}F_{m}\left(\begin{array}[c]{ccc}a_{1}&\ldots&a_{n}\\ b_{1}&\ldots&b_{m}\end{array};x\right)=$
(3.29) $(b_{1}-1)\sum_{k\geq 0}\frac{(-1)^{k}}{b_{1}-1+k}\frac{x^{k}}{k!}\frac{\prod_{s=1}^{n}\left(a_{s}\right)^{\left(k\right)}}{k!\prod_{s=2}^{m}\left(b_{s}\right)^{\left(k\right)}}~_{n}F_{m}\left(\begin{array}[c]{ccc}a_{1}+k&\ldots&a_{n}+k\\ 1+k&\ldots&b_{m}+k\end{array};x\right)$
Proof.
i) We will utilize (2.12), (2.8) and (2.16). We get then
$_{n}F_{m}\left(\begin{array}[c]{ccc}a_{1}&\ldots&a_{n}\\ b_{1}&\ldots&b_{m}\end{array};(1-x)/2\right)=\sum_{j\geq 0}\frac{\prod_{s=1}^{n}\left(a_{s}\right)^{\left(j\right)}}{\prod_{s=1}^{m}\left(b_{s}\right)^{\left(j\right)}}\frac{(1-x)^{j}}{j!2^{j}}$
$=\sum_{j\geq 0}(-1)^{j}\frac{\prod_{s=1}^{n}\left(a_{s}\right)^{\left(j\right)}}{\prod_{s=1}^{m}\left(b_{s}\right)^{\left(j\right)}}\frac{1}{j!}\sum_{k=0}^{j}\tilde{e}_{j,k}(d,f)J_{k}(x|d,f)$
$=\sum_{k\geq 0}\left(-1\right)^{k}J_{k}(x|d,f)\sum_{j\geq k}(-1\sum)^{j-k}\frac{\prod_{s=1}^{n}\left(a_{s}\right)^{\left(j\right)}}{\prod_{s=1}^{m}\left(b_{s}\right)^{\left(j\right)}}$
$\times\frac{1}{j!}\frac{(-1)^{j-k}j!\left(f+k\right)^{(j-k)}}{\left(j-k\right)!\left(d+f+k-1\right)^{\left(k\right)}\left(d+f+2k\right)^{\left(j-k\right)}}$
$=\sum_{k\geq 0}J_{k}(-x|f,d)\frac{\prod_{s=1}^{n}\left(a_{s}\right)^{\left(k\right)}}{\prod_{s=1}^{m}\left(b_{s}\right)^{\left(k\right)}\left(d+f+k-1\right)^{\left(k\right)}}$
$\times\sum_{j\geq k}\frac{\prod_{s=1}^{n}\left(a_{s}+k\right)^{\left(j-k\right)}}{\prod_{s=1}^{m}\left(b_{s}+k\right)^{\left(j-k\right)}}\frac{\left(f+k\right)^{(j-k)}}{\left(j-k\right)!\left(d+f+2k\right)^{\left(j-k\right)}}$
$=\sum_{k\geq 0}J_{k}(-x|f,d)\frac{\prod_{s=1}^{n}\left(a_{s}\right)^{\left(k\right)}}{\prod_{s=1}^{m}\left(b_{s}\right)^{\left(k\right)}\left(d+f+k-1\right)^{\left(k\right)}}$
$\times~_{n+1}F_{m+1}\left(\begin{array}[c]{cccc}a_{1+}+k&\ldots&a_{n}+k&f+k\\ b_{1}+k&\ldots&b_{m}+k&\left(d+f+2k\right)\end{array};1\right)$
ii) We proceed in the similar way by this time we will use polynomials $K_{j}$, hence we will use (2.13) and (2.18) and in the last stage we change 1x1-x to $x$.
$_{n}F_{m}\left(\begin{array}[c]{ccc}a_{1}&\ldots&a_{n}\\ b_{1}&\ldots&b_{m}\end{array};(1-x)\right)=\sum_{k\geq 0}(-1)^{k}\frac{\prod_{s=1}^{n}\left(a_{s}\right)^{\left(k\right)}}{\prod_{s=1}^{m}\left(b_{s}\right)^{\left(k\right)}}\frac{(x-1)^{k}}{k!}$
$=\sum_{k\geq 0}(-1)^{k}\frac{\prod_{s=1}^{n}\left(a_{s}\right)^{\left(k\right)}}{\prod_{s=1}^{m}\left(b_{s}\right)^{\left(k\right)}}\frac{1}{k!}\sum_{j=0}^{k}\tilde{e}_{k,j}(d,f)K_{j}(x|d,f)$
$=\sum_{j\geq 0}(-1)^{j}K_{j}(x|d,f)\sum_{k\geq j}\frac{(-1)^{k-j}}{k!}$
$\times\frac{\prod_{s=1}^{n}\left(a_{s}\right)^{\left(k\right)}}{\prod_{s=1}^{m}\left(b_{s}\right)^{\left(k\right)}}\frac{(-1)^{k-j}k!\left(b+j\right)^{\left(k-m\right)}}{(k-j)!\left(a+b+m-1\right)^{\left(j\right)}\left(a+b+2m\right)^{\left(k-j\right)}}.$
iii) We use (2.1) and get
$_{n}F_{m}\left(\begin{array}[c]{ccc}a_{1}&\ldots&a_{n}\\ b_{1}&\ldots&b_{m}\end{array};x\right)=(b_{1}-1)\sum_{j\geq 0}\frac{\prod_{s=1}^{n}\left(a_{s}\right)^{\left(j\right)}}{\prod_{s=2}^{m}\left(b_{s}\right)^{\left(j\right)}}\frac{x^{j}}{j!}\frac{1}{j!}\sum_{k=0}^{j}\binom{j}{k}\frac{(-1)^{k}}{b_{1}-1+k}$
$=(b_{1}-1)\sum_{k\geq 0}\frac{(-1)^{k}}{b_{1}-1+k}\frac{x^{k}}{k!}\sum_{j\geq k}\frac{x^{j-k}}{(j-k)!}\frac{\prod_{s=1}^{n}\left(a_{s}\right)^{\left(j\right)}}{(1)^{\left(j\right)}\prod_{s=2}^{m}\left(b_{s}\right)^{\left(j\right)}}$
$=(b_{1}-1)\sum_{k\geq 0}\frac{(-1)^{k}}{b_{1}-1+k}\frac{x^{k}}{k!}\frac{\prod_{s=1}^{n}\left(a_{s}\right)^{\left(k\right)}}{\prod_{s=2}^{m}\left(b_{s}\right)^{\left(k\right)}}\sum_{j\geq k}\frac{x^{j-k}}{(j-k)!}\frac{\prod_{s=1}^{n}\left(a_{s}+k\right)^{\left(j-k\right)}}{(k+1)^{\left(j-k\right)}\prod_{s=2}^{m}\left(b_{s}+k\right)^{\left(j-k\right)}}$
$=(b_{1}-1)\sum_{k\geq 0}\frac{(-1)^{k}}{b_{1}-1+k}\frac{x^{k}}{k!}\frac{\prod_{s=1}^{n}\left(a_{s}\right)^{\left(k\right)}}{k!\prod_{s=2}^{m}\left(b_{s}\right)^{\left(k\right)}}~_{n}F_{m}\left(\begin{array}[c]{ccc}a_{1}+k&\ldots&a_{n}+k\\ 1+k&\ldots&b_{m}+k\end{array};x\right)$
Corollary 2 (Hypergeometric case).
@@
(3.33) $_{3}F_{2}\left(\begin{array}[c]{ccc}a+j&b+j&f+j\\ c+j&d+f+2j&\end{array};1\right)=\frac{\Gamma(d+f)\Gamma(c+d-a-b)}{\Gamma(d+f-a)\Gamma(c+d-b)}\times$
$\sum_{j\geq 0}J_{j}\left(x|f,d\right)\frac{\left(a\right)^{\left(j\right)}\left(b\right)^{\left(j\right)}\left(d+f\right)^{\left(2j\right)}}{\left(c\right)^{\left(j\right)}\left(d+f+j-1\right)^{\left(j\right)}\left(d+f-a\right)^{\left(j\right)}\left(c+d-b\right)^{\left(j\right)}}$
(3.36) $\times~_{3}F_{2}\left(\begin{array}[c]{ccc}a+j&c-b&c-f\\ c+j&c+d-b+j&\end{array};1\right).$
@@
(3.40) $\times~_{3}F_{2}\left(\begin{array}[c]{ccc}a+j&b+j&f+j\\ c+j&d+f+2j&\end{array};1\right)=\frac{\Gamma(d+f)\Gamma(c+d-a-b)}{\Gamma(d+f-a)\Gamma(c+d-b)}$
$\sum_{j\geq 0}K_{j}\left(x|f,d\right)\frac{\left(a\right)^{\left(j\right)}\left(b\right)^{\left(j\right)}\left(d+f\right)^{\left(2j\right)}}{\left(c\right)^{\left(j\right)}\left(d+f+j-1\right)^{\left(j\right)}\left(d+f-a\right)^{\left(j\right)}\left(c+d-b\right)^{\left(j\right)}}$
(3.43) $\times~_{3}F_{2}\left(\begin{array}[c]{ccc}a+j&c-b&c-f\\ c+j&c+d-b+j&\end{array};1\right).$
@@
(3.44) $_{2}F_{1}\left(a,b;c;x\right)=\left(c-1\right)\sum_{j\geq 0}\frac{(-1)^{j}\left(a\right)^{\left(j\right)}\left(b\right)^{\left(j\right)}x^{j}}{\left(c-1+j\right)\left(j!\right)^{2}}~_{2}F_{1}\left(a+j,b+j;j+1;x\right)$
$=(1-x)^{1-a-b}\left(c-1\right)$